# X. Wang1,2,+    Siming Fu1,+    Qihan Huang1,2    Wanggui He1    Hao Jiang1,† 1

### v1

@@ 3.4 Multi-subject Cross-attention
$\hat{\mathbf{z}}_{img}=\text{Softmax}\left(\frac{\mathbf{Q}\mathbf{K}_{i}}{\sqrt{d}}+\mathbf{M}\right)\mathbf{V}_{i}$ (5)
Herein, $\mathbf{M}$ represents the amalgamation of all subject-specific masks, $\text{Concat}(\mathbf{M}_{0},\ldots,\mathbf{M}_{n})$.
However, an inherent limitation arises when a query patch token is ubiquitously masked across all referenced subjects or remains unmasked (in instances of overlapping bounding boxes), thereby diminishing the intended efficacy of multi-subject cross-attention. To counteract this, we introduce dummy tokens initialized at random preceding the image tokens to symbolize the background. This strategy is instrumental in ensuring that text conditions predominantly govern areas devoid of any guided layout, thereby considered irrelevant to the referenced images. Following the acquisition of $\mathbf{A}$, we apply $\mathbf{M}_{bg}$ to seamlessly mask these tokens within $\hat{\mathbf{z}}_{img}$, as illustrated:
@@
where $\mathbf{M}_{bg}$ is articulated as a binary mask, with elements within the subject bounding boxes designated as zero. Notably, certain studies have sought to resolve these conflicts through the application of objectives on attention maps within the cross-attention mechanism. A series of rigorous experiments have been conducted to substantiate our design, with the details elucidated in Section 4.4.
@@ 4 Experiments
Table 2: Quantitative comparison on MS-Diffusion and baselines. Bold and underline represent the highest and second-highest metrics, respectively. For single-subject, the results for IP-Adapter, Emu2, and Kosmos-G were obtained from [20], while the rest are reported in the corresponding papers. * denotes the model is fine-tuned on DreamBench.
@@
4.1 Experiment Setup
Datasets.
For training, we utilize an in-house video dataset that contains 3.6M video clips. For evaluation, we measure the single-subject and multi-subject performance on DreamBench [28] and MS-Bench, respectively. DreamBench includes 30 subjects and 25 prompts and we preset all input boxes to [0.25, 0.25, 0.75, 0.75]. To thoroughly assess the performance of multi-subject personalization, we have established a new evaluation standard, MS-Bench, which includes 40 subjects and 1148 combinations, each combination having up to 6 prompts, totaling 4488 distinct test samples. Details of datasets are provided in Section A and Section B.
Evaluation metrics.
Following previous works, we measure the performance through image and text fidelity. To assess image fidelity, we employ cosine similarity measures between generated images and subject images within CLIP [25] and DINO [4] spaces, referred to as CLIP-I and DINO, respectively. For text fidelity, we utilize cosine similarity between generated images and text prompts in CLIP space, denoted as CLIP-T. In multi-subject personalization, using the average fidelity to reflect image fidelity is insufficient, as it fails to reveal cases of subject neglect. We further employ the product of multi-subject DINO, denoted as M-DINO, to indicate whether each subject has been recreated in the results.
Baselines.
For single-subject personalization, we compare our model with methods mentioned in Table 1. Emu2 [30], Kosmos-G [19], and $\lambda$-ECLIPSE [20] are all MLLM-based methods, while $\lambda$-ECLIPSE is reported to have better performance. Therefore, we select SSR-Encoder [40] and $\lambda$-ECLIPSE [20] as baselines for multi-subject personalization. The implementation details of MS-Diffusion and these methods are contained in Section C.
@@ 4.2 Single-subject Comparison
In the single-subject comparison, a detailed examination is carried out utilizing both qualitative and quantitative comparisons to gauge the performance of different methodologies. On the qualitative front, as shown in Figure 4, MS-Diffusion shows an exceptional ability to generate single-subject images with high fidelity and detail retention. In quantitative results provided in Table 2, MS-Diffusion also achieves competitive scores, with obviously the highest DINO and CLIP-T scores at 0.671 and 0.321 respectively on zero-shot scenarios, and leading CLIP-I score of 0.792. For tuning methods, MS-Diffusion outperforms baselines in all metrics. As discussed in [28], DINO more accurately captures the similarity in details between the results and the labels, whereas CLIP-I may exhibit high scores in situations of background overfitting, resulting in a clear advantage for DINO, but a slight disadvantage for CLIP-I of MS-Diffusion.
@@ 4.4 Ablation Study
We conduct an ablation experiment on the proposed two modules, grounding resampler (GRS) and multi-subject cross-attention (MCA), to validate their effects. For GRS, we replace it with a linear projection layer and a normal resampler [1, 38]. Results in Table 3 indicate that the resampler-like image projector significantly enhances the details, as evidenced by DINO being obviouly higher than the linear projector. Moreover, The substantial improvement in multi-object image fidelity by GRS reflects the critical role of the information carried by grounding tokens in multi-object generation. As a key module for resolving conflicts, removing MCA results in a noticeable degradation of text fidelity, especially in multi-subject generation. The combined use of both modules ensures that MS-Diffusion maintains high image and text fidelity simultaneously.
@@
As mentioned in Section 3.4, we have explored the indispensable role of explicit layout guidance (LG), including grounding tokens and MCA. A straightforward approach to implicitly utilizing layout involves incorporating an attention loss during training. Besides the image attention loss (IAL), we also introduce text attention loss (TAL) to training by setting the original cross-attention layers trainable. The detailed loss definition is provided in Section F. As illustrated in Table 3, an objective to guide the image cross-attention helps the personalization hardly at all. TAL has somewhat resolved the conflict issues, but its performance is inferior to MS-Diffusion while introducing additional training parameters. We consider the inclusion of LG necessary and rational, not merely for the performance enhancements it offers, but also because it effectively resolves the various multi-object generation issues highlighted in Figure 2.
@@ Appendix A Training Dataset Construction Pipeline
Figure 6 illustrates the construction pipeline of the training dataset. Firstly, we randomly select two frames from a video clip, one as the reference and the other as the ground truth. Both frames are captioned by BLIP-2 [16]. Secondly, we utilize a NER model [^1] https://spacy.io/ to extract entities from the caption. Entities and images are then input into Grounding DINO [18] to obtain the boxes, which are parts of the final input of the model. Taking the boxes as prompts of SAM [13], we can further obtain segmentation masks to extract subjects from the reference image. Since the entities in different frames can be different, we design a subject matcher, which finds the correspondence between the frames by conducting Hungarian Algorithm on the entity image embeddings. Frames in a video typically contain the same entities but exhibit clear differences in details such as angles and poses. This makes them highly suitable as training data for personalized image generation models, which can help mitigate the model’s tendency to copy-and-paste.
@@ Appendix B Details of MS-Bench
Table 4: Explanation of MS-Bench. Each combination type has preset prompts and boxes. [S] represents prompt variations about the scene, including "in a room", "in the jungle", "in the snow", "on the beach", "on the grass", and "on a cobblestone street".
@@
To construct MS-Bench, we collect subjects from [28, 9, 14], the Internet [^2] https://unsplash.com/, and an internal dataset that does not overlap with the training set. MS-Bench contains four data types and 13 combination types with two or three subjects. We provide the details in Table 4. Each combination type other than those related to the scene has 6 prompt variations. There are 1148 combinations and 4488 evaluation samples, where entities and boxes are subject categories and preset layouts. Compared to other multi-subject benchmarks, our MS-Bench ensures that the model performance can be reflected comprehensively in abundant cases.
@@ Appendix C Experiment Settings
The pre-trained model employed in MS-Diffusion is Stable Diffusion XL (SDXL) [23]. Implemented by Pytorch 2.0.1 and Diffusers 0.23.1, our model is trained on 16 A100 GPUs for 120k steps with a batch size of 8 and a learning rate of 1e-4. Following the training of [38], we set $\gamma=1.0$ in cross-attention layers and dropped the text and image condition using the same probability. To ensure the model is not dependent on the grounding tokens (Section 3.3), we also randomly drop them with a probability of 0.1. We generate five images for each sample during the inference, with unconditional guidance scale and $\gamma$ set to 7.5 and 0.6, respectively, to get better results.
@@
  • BLIP-Diffusion [15] utilizes BLIP-2 [16] to unite the text and image embeddings. We implement it in the qualitative comparison using Diffusers.
  • IP-Adapter [38] also uses image prompt as the condition. We run qualitative samples on their official code with the scale set to 0.5 recommended by the paper. Considering fairness, we use the result of SDXL in Table 2.
  • SSR-Encoder [40] design a query network to extract the specified subject of a single image, which enables it to finish multi-subject generation. We leverage it as one of the baselines in multi-subject personalization. For performance comparison, we employed the official code provided, alongside the default hyperparameters specified in the code repository.
  • $\lambda$-ECLIPSE [20] trains an independent multi-modal encoder and employs Kandinsky as the generative backbone. Since it outperforms other MLLM-based approaches, we choose it to be the representative. To facilitate a comparative analysis of performance, the study utilized the officially provided code, in conjunction with the default hyperparameters delineated within the corresponding code repository.
@@ Appendix E More Results of Multi-subject Personalization
We provide additional multi-subject personalized images based on MS-Bench in Figure LABEL:fig:multi_app. The results encompass various combination types, fully demonstrating the generalizability and robustness of MS-Diffusion. When the scene changes freely according to the text, the details of the subject are preserved without being affected. In addition to common parallel combinations, MS-Diffusion also performs well in personalized generation for combinations with certain overlapping areas, such as "living+midwearing" and "object+scene".
Appendix F Layout Guidance
Cross-attention maps can intuitively reflect the condition-image attribution relation [31, 11]. Recent works [35, 37] have studied utilizing an objective on the cross-attention maps in multi-subject generation. The objective exists only in training, considered implicit and insufficient to handle multi-subject conflicts. We have provided the performance comparison between our explicit layout guidance and attention loss in Section 4.4. For a single cross-attention layer, the attention loss of the $j$th subject is calculated by:
@@ Appendix F Layout Guidance
To further explore the layout control capability of MS-Diffusion, we provide qualitative results in Figure LABEL:fig:layout. It can be demonstrated that MS-Diffusion can generate images that adhere to layout conditions, even in the case of two instances of the same category. However, the generated positions are not entirely accurate, especially in "a cat and a cat on the grass", illustrating that the layout condition is relatively weak compared to text and image prompts in the personalization task.
Appendix G Integration with ControlNet
@@
Appendix H Text-Image Attribution Analysis
In MS-Diffusion, our focus is primarily on resolving conflicts between subjects without altering the control mechanism of the text. While some approaches [35, 34, 42, 12] in non-personalized text-to-image tasks address multi-object generation conflicts through text cross-attention, this inevitably requires tuning the diffusion model’s parameters, thereby affecting the plug-and-play nature, which is not preferred by us. As demonstrated in the text-image attribution analysis presented in Figure 7, the control of multiple objects by text in our model is also quite evident. This may be related to the explicit layout guidance for subjects, since the images and text condition jointly in the generation process. We also attempted to control text cross-attention using the same mechanism as in Section 3.4, but no differences were observed in the results.
@@ Appendix H Text-Image Attribution Analysis
Appendix I Subject Interpolation

### v2

@@ 3.5 Multi-subject Cross-attention
$\hat{\mathbf{z}}_{img}=\text{Softmax}\left(\frac{\mathbf{Q}{\mathbf{K}^{\top}_{i}}}{\sqrt{d}}+\mathbf{M}\right)\mathbf{V}_{i}$ (5)
Herein, $\mathbf{M}$ represents the amalgamation of all subject-specific masks, $\text{Concat}(\mathbf{M}_{0},\ldots,\mathbf{M}_{n})$. In this way, the model ensures each subject to be represented in a certain area, thus resolving the issues of subject neglect and conflict in Figure 2.
However, an inherent limitation arises when a query patch token is ubiquitously masked across all referenced subjects or remains unmasked (in instances of overlapping bounding boxes), thereby diminishing the intended efficacy of multi-subject cross-attention. To counteract this, we introduce dummy tokens initialized at random preceding the image tokens to symbolize the background. This strategy is instrumental in ensuring that text conditions predominantly govern areas devoid of any guided layout, thereby solving the subject overcontrol issue in Figure 2. Following the acquisition of $\mathbf{A}$, we apply $\mathbf{M}_{bg}$ to seamlessly mask these tokens within $\hat{\mathbf{z}}_{img}$, as illustrated:
@@
where $\mathbf{M}_{bg}$ is articulated as a binary mask, with elements within the subject bounding boxes designated as zero. In contrast to the methods discussed in Section 3.2, our proposed multi-subject cross-attention directly manages the image conditions by employing masked image cross-attention in the targeted areas. While addressing multi-object conflicts, our method ensures that text conditions remain unaffected, as evidenced by the significantly higher text adherence capability shown in Table 2. Notably, certain studies have sought to resolve these conflicts through the application of objectives on attention maps within the cross-attention mechanism. A series of rigorous experiments have been conducted to substantiate our design, with the details elucidated in Section 4.4.
@@ 4 Experiments
4.1 Experiment Setup
Datasets.
For training, we utilize an in-house video dataset that contains 3.6M video clips. For evaluation, we measure the single-subject and multi-subject performance on DreamBench (Ruiz et al., 2023) and MS-Bench, respectively. DreamBench includes 30 subjects and 25 prompts and we preset all input boxes to [0.25, 0.25, 0.75, 0.75]. To thoroughly assess the performance of multi-subject personalization, we have established a new evaluation standard, MS-Bench, which includes 40 subjects and 1148 combinations, each combination having up to 6 prompts, totaling 4488 distinct test samples. Details of datasets are provided in Section A and Section B.
Evaluation metrics.
Following previous works, we measure the performance through image and text fidelity. To assess image fidelity, we employ cosine similarity measures between generated images and subject images within CLIP (Radford et al., 2021) and DINO (Caron et al., 2021) spaces, referred to as CLIP-I and DINO, respectively. For text fidelity, we utilize cosine similarity between generated images and text prompts in CLIP space, denoted as CLIP-T. In multi-subject personalization, using the average fidelity to reflect image fidelity is insufficient, as it fails to reveal cases of subject neglect. We further employ the product of multi-subject DINO, denoted as M-DINO, to indicate whether each subject has been recreated in the results.
Baselines.
For single-subject personalization, we compare our model with methods mentioned in Table 1. Emu2 (Sun et al., 2023), Kosmos-G (Pan et al., 2023), and $\lambda$-ECLIPSE (Patel et al., 2024) are all MLLM-based methods, while $\lambda$-ECLIPSE is reported to have better performance. Therefore, we select SSR-Encoder (Zhang et al., 2024) and $\lambda$-ECLIPSE (Patel et al., 2024) as baselines for multi-subject personalization. Considering the fairness, the qualitative results of MS-Diffusion are generated without any fine-tuning on benchmarks. The implementation details of MS-Diffusion and these methods are contained in Section C.
@@ 4.1 Experiment Setup
Table 2: Quantitative comparison on MS-Diffusion and baselines. Bold and underline represent the highest and second-highest metrics, respectively. For single-subject, the results for IP-Adapter, Emu2, and Kosmos-G were obtained from $\lambda$-ECLIPSE (Patel et al., 2024), while the rest are reported in the corresponding papers. * denotes the model is fine-tuned on DreamBench.
@@ 4.2 Single-subject Comparison
In the single-subject comparison, a detailed examination is carried out utilizing both qualitative and quantitative comparisons to gauge the performance of different methodologies. On the qualitative front, as shown in Figure 4, MS-Diffusion shows an exceptional ability to generate single-subject images with high fidelity and detail retention. In quantitative results provided in Table 2, MS-Diffusion also achieves competitive scores, with obviously the highest DINO and CLIP-T scores at 0.671 and 0.321 respectively on zero-shot scenarios, and leading CLIP-I score of 0.792. For tuning methods, MS-Diffusion outperforms baselines in all metrics. As discussed in DreamBooth (Ruiz et al., 2023), DINO more accurately captures the similarity in details between the results and the labels, whereas CLIP-I may exhibit high scores in situations of background overfitting, resulting in a clear advantage for DINO, but a slight disadvantage for CLIP-I of MS-Diffusion.
@@ 4.3 Multi-subject Comparison
Figure 6: Visualization results of module ablation. Models without RS or GRS have an obvious decrease in the detail-preserving capability. For the multi-subject generation, the model without MCA cannot handle the subject conflicts.
@@ 4.4 Ablation Study
We conduct an ablation experiment on the proposed two modules, grounding resampler (GRS) and multi-subject cross-attention (MCA), to validate their effects. For GRS, we replace it with a linear projection layer and a normal resampler (Alayrac et al., 2022; Ye et al., 2023). Results in Table 3 indicate that the resampler-like image projector significantly enhances the details, as evidenced by DINO being obviouly higher than the linear projector. Moreover, The substantial improvement in multi-object image fidelity by GRS reflects the critical role of the information carried by grounding tokens in multi-object generation. As a key module for resolving conflicts, removing MCA results in a noticeable degradation of text fidelity, especially in multi-subject generation. The combined use of both modules ensures that MS-Diffusion maintains high image and text fidelity simultaneously. We provide visualization results regarding the module ablation in Figure 6. As clearly reflected by the qualitative examples, GRS enhances the details and MCA handles the conflicts.
@@
As mentioned in Section 3.5, we have explored the indispensable role of explicit layout guidance (LG), including grounding tokens and MCA. A straightforward approach to implicitly utilizing layout involves incorporating an attention loss during training. Besides the image attention loss (IAL), we also introduce text attention loss (TAL) to training by setting the original cross-attention layers trainable. The detailed loss definition is provided in Section G. As illustrated in Table 3, an objective to guide the image cross-attention helps the personalization hardly at all. TAL has somewhat resolved the conflict issues, but its performance is inferior to MS-Diffusion while introducing additional training parameters. We consider the inclusion of LG necessary and rational, not merely for the performance enhancements it offers, but also because it effectively resolves the various multi-object generation issues highlighted in Figure 2.
@@ Appendix A Training Dataset Construction Pipeline
Figure 7 illustrates the construction pipeline of the training dataset. Firstly, we randomly select two frames from a video clip, one as the reference and the other as the ground truth. Both frames are captioned by BLIP-2 (Li et al., 2023b). Secondly, we utilize a NER model [^1] https://spacy.io/ to extract entities from the caption. Entities and images are then input into Grounding DINO (Liu et al., 2023a) to obtain the boxes, which are parts of the final input of the model. Taking the boxes as prompts of SAM (Kirillov et al., 2023), we can further obtain segmentation masks to extract subjects from the reference image. Since the entities in different frames can be different, we design a subject matcher, which finds the correspondence between the frames by conducting Hungarian Algorithm on the entity image embeddings. Frames in a video typically contain the same entities but exhibit clear differences in details such as angles and poses. This makes them highly suitable as training data for personalized image generation models, which can help mitigate the model’s tendency to copy-and-paste.
@@ Appendix B Details of MS-Bench
Table 4: Explanation of MS-Bench. Each combination type has preset prompts and boxes. [S] represents prompt variations about the scene, including ”in a room”, ”in the jungle”, ”in the snow”, ”on the beach”, ”on the grass”, and ”on a cobblestone street”.
@@
To construct MS-Bench, we collect subjects from previous studies (Ruiz et al., 2023; Gal et al., 2023; Kumari et al., 2023), the Internet [^2] https://unsplash.com/, and an internal dataset that does not overlap with the training set. MS-Bench contains four data types and 13 combination types with two or three subjects. We provide the details in Table 4. Each combination type other than those related to the scene has 6 prompt variations. There are 1148 combinations and 4488 evaluation samples, where entities and boxes are subject categories and preset layouts. Compared to other multi-subject benchmarks, our MS-Bench ensures that the model performance can be reflected comprehensively in abundant cases.
@@ Appendix C Experiment Settings
The pre-trained model employed in MS-Diffusion is Stable Diffusion XL (SDXL) (Podell et al., 2023). Implemented by Pytorch 2.0.1 and Diffusers 0.23.1, our model is trained on 16 A100 GPUs for 120k steps with a batch size of 8 and a learning rate of 1e-4. Following the training of IP-adapter (Ye et al., 2023), we set $\gamma=1.0$ in cross-attention layers and dropped the text and image condition using the same probability. To ensure the model is not dependent on the grounding tokens (Section 3.4), we also randomly drop them with a probability of 0.1. We generate five images for each sample during the inference, with unconditional guidance scale and $\gamma$ set to 7.5 and 0.6, respectively, to get better results.
@@
  • BLIP-Diffusion (Li et al., 2023a) utilizes BLIP-2 (Li et al., 2023b) to unite the text and image embeddings. We implement it in the qualitative comparison using Diffusers.
  • IP-Adapter (Ye et al., 2023) also uses image prompt as the condition. We run qualitative samples on their official code with the scale set to 0.5 recommended by the paper. Considering fairness, we use the result of SDXL in Table 2.
  • SSR-Encoder (Zhang et al., 2024) design a query network to extract the specified subject of a single image, which enables it to finish multi-subject generation. We leverage it as one of the baselines in multi-subject personalization. For performance comparison, we employed the official code provided, alongside the default hyperparameters specified in the code repository.
  • $\lambda$-ECLIPSE (Patel et al., 2024) trains an independent multi-modal encoder and employs Kandinsky as the generative backbone. Since it outperforms other MLLM-based approaches, we choose it to be the representative. To facilitate a comparative analysis of performance, the study utilized the officially provided code, in conjunction with the default hyperparameters delineated within the corresponding code repository.
@@ Appendix D More Results of Single-subject Personalization
One of the limitations in detail preservation is the insufficient capability of the image encoder. We provide some uncommon examples in Figure LABEL:fig:uncommon. MS-Diffusion utilizes the CLIP image encoder, which results in the loss of some details in uncommon and complex cases. However, it still significantly outperforms state-of-the-art methods benefiting from the proposed grounding resampler.
@@ Appendix E More Results of Multi-subject Personalization
We provide additional multi-subject personalized images based on MS-Bench in Figure LABEL:fig:multi_app. The results encompass various combination types, fully demonstrating the generalizability and robustness of MS-Diffusion. When the scene changes freely according to the text, the details of the subject are preserved without being affected. In addition to common parallel combinations, MS-Diffusion also performs well in personalized generation for combinations with certain overlapping areas, such as ”living+midwearing” and ”object+scene”.
Figure 8: Qualitative comparison of Mix-of-Show, Cones2, and MS-Diffusion. MS-Diffusion outperforms Mix-of-Show and Cones2 in details preserving in the first row. Mix-of-Show and Cones2 struggle in handling the interactions between subjects.
Figure 9: Qualitative comparison of Break-A-Scene and MS-Diffusion. MS-Diffusion gets comparable results by extracting subjects from a single image. Break-A-Scene tends to overfit the input image (the bird pose in the first row and the white creature sitting on the black bowl in the second row).
Appendix F Comparison with Tuning-based Methods
While zero-shot methods like MS-Diffusion can decrease the tuning cost, they may also suffer from performance degradation compared to tuning-based approaches due to the limitations of pre-training scale. However, MS-Diffusion indicates comparable performance in single-subject quantitative results in Table 2. Since the proposed MS-Bench can be too large for tuning-based methods, we provide qualitative comparison results with Mix-of-Show (Gu et al., 2023), Cones2 (Liu et al., 2023b), and Break-A-Scene (Avrahami et al., 2023) in Figure 8 and Figure 9. The results show that MS-Diffusion achieves comparable results to tuning-based methods. Mix-of-Show and Cones2 face certain issues when handling multiple subjects with complex interactions (e.g., the example of a dog wearing a dress). Break-A-Scene tends to overfit the original image’s interactions (e.g., the pose of birds and the white creature sitting on the black bowl). While avoiding these issues, MS-Diffusion requires no test-time tuning and only one subject image during inference, unlike the tuning-based methods that need additional time and multiple images for tuning.
Appendix G Layout Guidance
Cross-attention maps can intuitively reflect the condition-image attribution relation (Tang et al., 2023; Hertz et al., 2023). Recent works (Wang et al., 2023; Wei et al., 2024) have studied utilizing an objective on the cross-attention maps in multi-subject generation. The objective exists only in training, considered implicit and insufficient to handle multi-subject conflicts. We have provided the performance comparison between our explicit layout guidance and attention loss in Section 4.4. For a single cross-attention layer, the attention loss of the $j$th subject is calculated by:
@@ Appendix G Layout Guidance
To further explore the layout control capability of MS-Diffusion, we provide qualitative results in Figure LABEL:fig:layout. It can be demonstrated that MS-Diffusion can generate images that adhere to layout conditions, even in the case of two instances of the same category. However, the generated positions are not entirely accurate, especially in ”a cat and a cat on the grass”, illustrating that the layout condition is relatively weak compared to text and image prompts in the personalization task.
Appendix H Integration with ControlNet
@@
Appendix I Multi-subject Interaction
Unlike image editing (Kawar et al., 2023; Brooks et al., 2023), personalized image generation features a high degree of freedom, enabling effective handling of interactions among different elements. Benefiting from its architecture design that does not impact the base model, MS-Diffusion successfully inherits the multi-subject interaction capabilities of the base model. As illustrated in Figure LABEL:fig:interaction, MS-Diffusion can flexibly manage interactions between reference subjects, even when there is overlap among these objects, thereby demonstrating its potential in practical applications.
Appendix J Human and Anime Personalization
Human and anime subjects are popular in the use of personalized model. We provide results of MS-Diffusion on human and anime subjects in Figure LABEL:fig:human. Some research (Wang et al., 2024a; Li et al., 2024) has explored human personalization in text-to-image diffusion models. By utilizing a face encoder (Deng et al., 2019) and training on the face dataset, MS-Diffusion can also be extended to a personalized model for these subjects.
Appendix K Text-Image Attribution Analysis
In MS-Diffusion, our focus is primarily on resolving conflicts between subjects without altering the control mechanism of the text. While some approaches (Wang et al., 2023; 2024b; Zhou et al., 2024; Kim et al., 2023) in non-personalized text-to-image tasks address multi-object generation conflicts through text cross-attention, this inevitably requires tuning the diffusion model’s parameters, thereby affecting the plug-and-play nature, which is not preferred by us. As demonstrated in the text-image attribution analysis presented in Figure 10, the control of multiple objects by text in our model is also quite evident. This may be related to the explicit layout guidance for subjects, since the images and text condition jointly in the generation process. We also attempted to control text cross-attention using the same mechanism as in Section 3.5, but no differences were observed in the results.
Appendix L Subject Interpolation